Wehner Company is currently manufacturing Part ABS-43, producing 57,800 units annually. The part is used in the production of several products made by Wehner. The cost per unit for ABS-43 is as follows: Direct materials $47.50 Direct labor 10.75 Variable overhead 3.30 Fixed overhead 4.30 Total $65.85 Of the total fixed overhead assigned to ABS-43, $13,121 is direct fixed overhead (the annual lease cost of machinery used to manufacture Part ABS-43), and the remainder is common fixed overhead. An outside supplier has offered to sell the part to Wehner for $61.39. There is no alternative use for the facilities currently used to produce the part. No significant non-unit-based overhead costs are incurred.

. Should Wehner Company make or buy Part ABS-43?
Wehner should___________ the part. This will produce total cost savings of $.

2. What is the maximum amount per unit that Wehner would be willing to pay to an outside supplier? Round your answer to the nearest cent.
$ per unit
Answer #1

1. Should Wehner Company make or buy Part ABS-43?

Total Relevant cost of Manufacturing = (Direct material + Direct Labor + Variable Overhead)*No of Unit Produced + direct fixed overhead

Total Relevant cost of Manufacturing = (47.50+ 10.75+ 3.30)*57,800 + 13,121

Total Relevant cost of Manufacturing =3,557,590+13,121

Total Relevant cost of Manufacturing = $ 3,570,711

Total Cost of Buying = 57800*61.39

Total Cost of Buying = $ 3,548,342

Total cost savings if part is bought = 3570711 - 3548342

Total cost savings if part is bought = $ 22,369

Answer

Decision : Wehner should buy the part . This will produce total cost savings of $ 22,369

2 Ans

Maximum Amount that Wehner would be willing to pay to an outside supplier = Total Cost to make/ no of units

Total Cost to Make : 3570711

No of units : 57800

Maximum Amount that Wehner would be willing to pay to an outside supplier = 3570711/57800=61.7770 = 61.78

Answer = $ 61.78